# Adjust for Trust: Mitigating Trust-Induced Inappropriate Reliance on AI Assistance

### v1

@@ 4.2 Mitigating Over-reliance with Counter-Explanations
Figure 6 shows that providing counter-explanations adaptively mitigates over-reliance when user trust is high.
@@
Trust-Adaptive Counter-Explanations Help.
@@
Counter-explanations are effective at reducing over-reliance when user trust is high, across all task settings, with 10–23% reduction in Over-Reliance, 19–36% reduction in Total Inappropriate Reliance, and 8–20% improvement in Final Decision Accuracy. When considering all interactions where the user and AI have different prediction, improvements are also observed in almost all cases but to a lesser extent.
@@
Persistent Counter-Explanations Are Not As Helpful.
@@
The effects become less pronounced in the Intervention Always condition. In the ArcO setting, users perform uniformly worse than in the No Intervention condition when trust is high. Users may be less inclined to closely evaluate counter-explanations when they are always shown.
4.3 Mitigating Under- and Over-Reliance Simultaneously
@@
We now investigate whether showing supporting explanations when trust is low ($<5$) and counter-explanations when trust is high ($>8$) in the same user-AI study yields complementary improvements in decision accuracy and inappropriate reliance.
@@
Figure 7 shows that providing different types of explanations based on user trust yields complementary performance improvements on both the ARC and Diagnosis tasks. The benefits observed by using supporting explanations during low trust mirror those previously observed in §4.1, while the benefits of using counter-explanation are similar to those observed in §4.2.
4.4 Intervention through Deceleration
Figure 8: Effect of decelerating interventions on reducing inappropriate reliance and improving decision accuracy.
@@ 4.4 Intervention through Deceleration
We have demonstrated that supporting and counter-explanations are useful for mitigating under- and over-reliance, respectively. We now investigate the utility of another type of intervention: slowing down the interaction to promote deliberation, without providing additional information to the user.
To mitigate under-reliance by decelerating the interaction, we display a “The AI is thinking…” message for 10 seconds before the AI prediction is revealed to the user. To mitigate over-reliance, we reveal the AI prediction and ask the user to carefully consider the AI advice; the user is made to wait 10 seconds before making their final decision.
Findings.
@@
Figure 8 shows the effect of the above decelerating interventions at mitigating inappropriate reliance in the ArcC task setting. We see that telling users that the AI is thinking is not particularly effective at reducing Total Inappropriate Reliance, even when trust is low. On the other hand, forcing users to consider the AI advice closely when trust is high improves Total Inappropriate Reliance and Final Decision Accuracy. Our findings extend those of Buçinca et al. (2021), showing that cognitive forcing is particularly useful for reducing over-reliance when user trust is high.
5 Discussion
Applying trust-adaptive interventions.
Can we model user trust?
In our setting, the AI assistant can observe the user’s trust level after each interaction. In appendix D, we provide an analysis of several heuristic-based and learning-based trust models on their ability to predict user trust levels based on user-AI interaction history instead. We find that these heuristics only achieve moderate correlation with user-reported trust levels, and are especially poor at predicting moments of high and low user trust. Our results point to the challenging nature of modeling user trust, and the inadequacy of surface-level interaction features. Instead, user-specific features such as users’ internal confidence, their prior experience with AI systems and task expertise may be better indicators of user trust.
Is all inappropriate reliance equally bad?
6 Conclusion and Future Directions
Our findings present an initial exploration into adapting AI behaviors based on user trust levels. We adopted a simple thresholding criterion for deciding when to intervene, but more sophisticated criteria that also account for user and AI confidence may have potential. Further, rather than looking at whether user trust is too high or low, we can consider whether the trust is calibrated with the assistant’s trustworthiness. High user trust may not be as undesirable when the AI is significantly more accurate than the user on the task. We hope our findings inspire the community to more closely consider the effect of user trust in user-AI interactions and the potential benefits of modeling and adapting to user trust levels.
Limitations
Our user studies were performed using a simulated AI assistant, rather than a real system such as a Large Language Model. Similar to Buçinca et al. (2021); Dhuliawala et al. (2023), we use a simulated AI for the controllability of AI accuracy and confidence calibration. However, a simulated AI may be unrealistic compared to real AI systems that people use. For instance, since the answer correctness is decided independently for each problem, our AI assistant may provide contradicting predictions for near-identical problems.
Furthermore, the behavior of Prolific participants interacting with an AI in a user study may differ from users of a live system in a real-world scenario McGrath (1995), particularly due to misalignment of motivations Deci et al. (1999). We test for the effect of two confounding variables: trust reporting and user experience with AI (Appendix E), finding that they do not affect user reliance behavior, but there may be other confounding variables we have not yet considered.
Finally, our findings would be more reliable if we could obtain data from more participants, more tasks and more interventions, however this is not possible due to financial restrictions and the dearth of professional doctors on the Prolific platform.
Ethical Considerations
Acknowledgments
We thank Rishi Vanukuru, Gregory Yauney, Sayan Ghosh and Jesse Zhang for their invaluable discussions and brainstorming, as well as Malihe Alikhani, Heather Culbertson, Suvodip Dey, Vardhan Dongre, Mert İnan, Sunnie Kim, Chaitanya Malaviya, Ana Marasović, Weiyan Shi, Anthony Sicilia, Swabha Swayamdipta, Dilek Hakkani-Tür and Gökhan Tür for their valuable feedback. We also thank Dhuliawala et al. (2023) for open-sourcing their user study interface, and in particular Vilém Zouhar for his assistance with setting up the user study and data logging. We are also very grateful to Aisha Baptiste for swiftly adding money to our Prolific account whenever required.
Tejas Srinivasan was supported by the Amazon ML Fellowship from the USC-Amazon Center for Secure and Trusted Machine Learning. This work was partially supported by a grant from the DARPA Friction and Accountability in Conversational Transactions (FACT) program. The opinions and findings expressed in this work do not necessarily reflect those of USC, Amazon, or DARPA.
Appendix A Task Details
ARC Task.
This task consists of answering science questions. Questions are sourced from the ARC dataset Clark et al. (2018), which consists of more than 7000 grade-school science multiple-choice questions written for examinations. The authors manually reviewed questions from this dataset and selected questions that were challenging (i.e. the correct answer is not immediately obvious, and at least one option was not obviously incorrect) but still understandable (did not contain any scientific jargon that laypeople may not be familiar with). All questions in the original dataset had four options, but the author only selected the correct answer and the most plausible incorrect option for the decision-making problem. The final filtered set consisted of 39 questions. For each of the 10 problem sequences that users solve, we sample 30 of the 39 questions without repetition.
Diagnosis Task.
This task consists of diagnosing patients based on patient symptoms. Patient symptoms are sourced from the DDXPlus dataset Fansi Tchango et al. (2022), which contains 1.3 million synthetic patients with a differential diagnosis. The symptoms are presented as either binary, categorical or continuous variables, but each symptom has a corresponding patient intake question and answer in English, which we translate into a descriptive third-person statement using GPT-4o.. For example, the intake question “Do you feel pain somewhere?” and patient response “Knee (R)” is translated into “The patient feels pain in their right knee.”. We filter down to questions with only 10–15 intake responses so that users do not need to spend a long time understanding the problem. To convert the task into a multiple-choice problem, we select the top three negative conditions from the differential diagnosis as the incorrect options. Our final set of problems includes 55 cases corresponding to eleven different conditions, which we use to sample 10 sequences of 30 problems each.
Appendix B User Study Details
Figure 10: Macro-aggregation results for mitigating under- and over-reliance. $n$ represents the number of users who have at least 3 interactions meeting the required criteria.
We present additional details about the user studies.
User Payment.
As mentioned earlier, users in the ARC task are paid a base payment of \$1.0, with an incentive of \$0.10 per correct answer. Users achieved 65–75% accuracy on the task, which translates to a bonus of \$$\approx\$2.0$ per user, or \$3.0 total payment. The tasks took a median time of 15 minutes to complete, which translates to a pay rate of \$12 per hour.
User Demographics.
Instructions.
Users were informed that they were participating in a research study. Figure 9 contains screenshots of the task instructions and payment details presented to users.
Appendix C Macro-Aggregation Results
In Figure 10, we observe trends similar to those observed in Section 4. However, when looking at the low/high trust subsets, most conditions have fewer than 10 users for calculating Under-Reliance and Over-Reliance, and fewer than 5 users for Total Inappropriate Reliance.
Appendix D Trust Modeling
@@
We experiment with several simple heuristics that are computed by looking at interaction outcomes. AIAcc5 estimates trust based on the AI’s accuracy in the last 5 interactions. CapabilityDiff calculates the difference between the accuracy of the AI’s prediction and the user’s initial decision, over all interactions so far. SmoothOutcomes updates the trust $\tau_{t}\in[-1,1]$ after interacton $t$, based on the AI’s prediction correctness $a_{t}\in\{0,1\}$
@@
where $r$ is a smoothing parameter. SmoothConfs is similar to the above, except the update term is weighted by the AI’s confidence $c^{AI}_{t}$
@@
Finally, we train a TrustModel, a linear regression model which estimates the change in user trust after each interaction.
@@
Table 1: Correlation of trust estimation methods with user-reported trust levels, and F1 for detecting moments of low and high trust.
@@
Table 1 shows correlation between trust scores estimated by the various heuristics and the user-reported trust levels, on both train and tests, along with the F1 for detecting moments of low and high user trust. We observe that all of the above methods achieve only moderate correlation with user-reported trust, and achieve very low F1 ($\leq 0.5$) at detecting both low and high trust instances. These results point to the challenging nature of modeling user trust, and the inadequacy of surface-level interaction features.
Appendix E Testing for Confounding Variables

### v2

@@ 7.2. Results
Figure 7 shows that providing counter-explanations only when user trust is high reduces over-reliance and improves users’ decision-making accuracy.
@@
Trust-adaptive counter-explanations reduce over-reliance (hypothesis H6).
@@
Counter-explanations are effective at reducing over-reliance when user trust is high, across all task settings, with 10–23% reduction in Over-Reliance, 19–36% reduction in Total Inappropriate Reliance, and 8–20% improvement in Final Decision Accuracy (all significant at $p<0.05$). When considering all interactions where the user and AI initially disagree, improvements are also observed in almost all cases but to a lesser extent.
@@
Persistent counter-explanations are not as helpful as trust-adaptive ones (hypothesis H7).
@@
The benefits of providing counter-explanations become less pronounced in the Intervention Always condition. In the ArcO setting, users perform uniformly worse than in the No Intervention condition when trust is high. These findings indicate that users may be less inclined to closely evaluate counter-explanations when they are always shown.
8. Experiment #3: Mitigating Under- and Over-Reliance Simultaneously
@@
Bringing these findings together, we investigate whether showing supporting explanations when trust is low (<5<5) and counter-explanations when trust is high (>8>8) in the same user-AI study yields complementary improvements in decision accuracy and inappropriate reliance.
  1. H8:
    Applying both interventions individually (showing supporting explanations when trust is low and counter-explanations when trust is high) will lead to lower Total Inappropriate Reliance and higher Final Decision Accuracy compared to applying each intervention individually.
8.1. Experiment Setup
8.1.1. Conditions
We compare the No Intervention baseline and the two earlier trust-adaptive intervention settings applied individually to the setting where the supporting explanation is shown when user trust is low and a counter-explanation is shown when trust is high. Due to limited budget for user studies, in this experiment we only study the two task settings corresponding to a calibrated AI assistant (ArcC and DiagC).
8.1.2. Evaluating Interventions
Since we are comparing interventions that mitigate both under-reliance and over-reliance, we primarily evaluate conditions based on their Total Inappropriate Relianceand users’ Final Decision Accuracy. We evaluate the four conditions across all interactions, regardless of trust level. To understand whether applying both interventions together yields the same results during moments of low and high trust as when they were applied individually, we also look at the subsets of interactions corresponding to low and high user trust.
Figure 8. Effect of providing supporting explanations (/) and counter-explanations (\), depending on user trust, within the same user session yields complementary benefits in inappropriate reliance and decision-making accuracy.
Effect of providing supporting explanations (/) and counter-explanations (\), depending on user trust, within the same user session
8.2. Results
@@
Figure 8 shows that providing different types of explanations based on user trust yields complementary performance improvements on both the ARC and Diagnosis tasks. When comparing conditions by considering all user-AI interactions in each, the condition where both types of interventions are applied depending on the trust level has the highest Final Decision Accuracy and lowest Total Inappropriate Reliance in both task settings. Further, the benefits observed by using supporting explanations during low trust mirror those observed in §6, while the benefits of counter-explanation are similar to those in §7. These findings indicate that when adapting AI behavior to simultaneously combat under- and over-reliance, the two types of interventions do not interfere with each other.
9. Experiment #4: Intervention through Deceleration
@@
We have demonstrated in Sections 6, 7, and 8 that supporting and counter-explanations are useful for mitigating under- and over-reliance, respectively. However, these need not be the only modes of AI behavior adaptation when user trust is low or high. We now investigate the utility of another type of intervention: slowing down the interaction to promote deliberation, without providing additional information to the user. Cognitive forcing functions (Park et al., 2019; Buçinca et al., 2021; Rastogi et al., 2022) have been widely studied as a tool to encourage users to deliberate and engage in slower, reflective System-2 thinking (Evans, 2003).
We experiment with two different decelerating interventions. To mitigate under-reliance by decelerating the interaction, we display a “The AI is thinking…” message for 10 seconds before the AI prediction is revealed to the user. We hypothesize that telling the users that the AI is reasoning more deeply (even though no additional reasoning is actually happening) could correct for users’ trust being miscalibrated with the AI assistant’s trustworthiness. To mitigate over-reliance, we reveal the AI prediction and ask the user to carefully consider the AI advice; the user must wait 10 seconds before being allowed to make their final decision. We hypothesize that these two decelerating interventions can mitigate inappropriate reliance by preventing users from making quick judgments based on their trust levels.
  1. H9:
    Informing users that the AI is doing additional reasoning when user trust is low can improve users’ Final Decision Accuracy and reduce Total Inappropriate Reliance.
  2. H10:
    When user trust is high, forcing users to pause before making a decision to rely on AI advice can reduce Total Inappropriate Reliance and improve Final Decision Accuracy.
9.1. Experiment Setup
9.1.1. Conditions
We experiment with the following conditions:
  1. (1)
    No Intervention: The AI’s behavior stays the same throughout the user study, regardless of users’ trust level.
  2. (2)
    Show “AI is Thinking…” when Trust is Low: Whenever the user’s trust level is below 5 out of 10, users are shown an “AI is thinking…” message for 10 seconds before the AI recommendation is revealed.
  3. (3)
    Add a Forced Pause when Trust is High: Whenever the user’s trust level is above 8 out of 10, after revealing the AI recommendation, users are forced to wait for 10 seconds before making their final decision.
  4. (4)
    Adaptively showing “AI is thinking” or adding forced pause: Similar to the intervention combination in §8, this condition is a combination of the previous two, where AI behavior is adapted when the user’s is either low or high, but the intervention to be applied depends on the specific trust level.
Figure 9. Effect of decelerating interventions on reducing inappropriate reliance and improving decision accuracy.
Effect of decelerating interventions on reducing inappropriate reliance
9.2. Results
Figure 9 shows the effect of the proposed interventions at mitigating inappropriate reliance in the ArcC task.
Showing users an “AI is thinking” message does not mitigate under-reliance (hypothesis H9)
We observe that telling users that the AI is thinking is not particularly effective at reducing Total Inappropriate Reliance, even when trust is low, and users’ Final Decision Accuracydoes not improve either.
Adding forced pauses reduces over-reliance (hypothesis H10)
@@ 9.2. Results
We find that forcing users to consider the AI advice closely when trust is high improves Total Inappropriate Reliance and Final Decision Accuracy. Our findings extend those of Buçinca et al. (2021), showing that cognitive forcing is particularly useful for reducing over-reliance when user trust is high.
10. Experiment #5: Trust-Adaptive Interventions with an LLM Assistant
Our studies so far have involved a simulated AI assistant where we control the assistant’s accuracy and calibration. We now investigate whether our findings generalize when users interact with an actual AI assistant, where we no longer have explicit control of model accuracy and confidence calibration. Specifically, we experiment with a Large Language Model (LLM) assistant. LLMs are now widely used to help people make better and more informed decisions, including with experts in high-stakes domains (Gumilar et al., 2024), but are prone to cause inappropriate reliance through the use of overconfident (Zhou et al., 2024) and sycophantic (Sharma et al., 2024) language.
  1. H11:
    If an LLM assistant has access to the user’s current trust level, adapting its behavior to provide either supporting or counter-explanations based on user trust level can improve users’ Final Decision Accuracy and reduce Total Inappropriate Reliance.
10.1. Experiment Setup
We compare two conditions: the Control condition where there is no intervention, and the Trust-Adaptive condition where the LLM assistance type depends on the user trust level. We recruited 30 participants for each of the two conditions, we were assisted by a LLaMA-3.1-8B (Grattafiori et al., 2024) model on the ARC decision-making task. [^7] Due to the sparsity of doctors on the Prolific platform, we found it difficult to conduct additional studies on the Diagnosis task, beyond those conducted in the previous experiments.
10.1.1. Generating LLM Recommendations
An LLM assistance includes a subset of the following: a prediction, a confidence estimate, a supporting explanation and a counter-explanation. We use a simple self-consistency framework (Wang et al., 2023) to generate all of the above in an efficient batched fashion during the user studies. We use chain-of-thought prompting (Wei et al., 2022) to generate 10 rationale-prediction pairs for each question. The majority predicted answer is selected as the LLM’s prediction, with its relative frequency representing the model confidence; we found self-consistency prompting to yield more calibrated estimates than prompting the LLM to output its confidence in its generation. In the Trust-Adaptive condition, if user trust is low, the LLM assistant selects one of the chain-of-thought rationales corresponding to the model prediction and provides it to the user as a supporting explanation. Similarly, when user trust is high, the LLM assistant would select a rationale corresponding to the other answer (if one exists among the 10 generated pairs) to show as a counter-explanation.
User Trust Metric No Intervention Trust-Adaptive LLM
<5<5 or >8>8 Decision Acc. 58.4% 67.8%
TotalInappRel. 83.5% 68.2%
<5<5 Under-Reliance 65.1% 44.4%
>8>8 Over-Reliance 56.4% 42.4%
Table 1. Effect of providing trust-adaptive supporting and counter-explanations when users are assisted by LLaMA-3.1 to answer questions from the ARC task.
10.2. Results
Table 1 compares the effect of both experiment conditions on user reliance outcomes. We observe that adapting LLM behavior to user trust improves users’ decision accuracy and reduces inappropriate reliance during moments of low and high trust, mirroring our findings in §8, thus validating hypothesis H11. Inspecting more closely, we see that under-reliance is reduced when user trust is low and over-reliance is reduced when trust is high. These findings indicate that simply having access to a trust signal can allow LLM assistants to improve human-LLM collaboration through trust-aware adaptation.
11. Analysis: Can We Model User Trust?
Our decision-making setting assumes that the AI assistant can observe the user’s trust level after each interaction. However, in real-world settings, the AI assistants may not have access to the trust signal. In the absence of this signal, it would be valuable for the AI assistant to be able to predict users’ trust levels based on the interaction history.
@@ 11. Analysis: Can We Model User Trust?
11.1. Trust Heuristics and Models
We experiment with several simple heuristics that are computed by looking at interaction outcomes
  1. (i)
    AIAcc5 estimates trust based on the AI’s accuracy in the last 5 interactions.
  2. (ii)
    CapabilityDiff calculates the difference between the accuracy of the AI’s prediction and the user’s initial decision, over all interactions so far.
  3. (iii)
    SmoothOutcomes updates trust $\tau_{t}\in[-1,1]$ after interaction $t$, based on AI prediction correctness $a_{t}\in\{0,1\}$
@@ 11.1. Trust Heuristics and Models
    where $r$ is a smoothing parameter.
  4. (iv)
    SmoothConfs is similar to the above, except the update term is weighted by the AI’s confidence $c^{AI}_{t}$
@@
  5. (v)
    We train a TrustModel, a linear regression model which estimates the change in user trust after each interaction.
@@
Table 2. Correlation of trust estimation methods with user-reported trust levels, and F1 for detecting moments of low and high trust.
11.2. Results
@@
Table 2 shows correlation between trust scores estimated by the various heuristics and the ground-truth user-reported trust levels, on both train and tests, along with the F1 for detecting moments of low and high user trust. We observe that all of the above methods achieve only moderate correlation with user-reported trust, and achieve very low F1 ($\leq 0.5$) at detecting both low and high trust instances. These results point to the challenging nature of modeling user trust, and the inadequacy of surface-level interaction features. Future work can model trust using more sophisticated interaction features (Fang et al., 2013).
12. Discussion
Implications for NLP systems.
While our findings highlight the benefits of adapting AI behavior to user trust and prior user-AI interactions, LLMs currently generate responses without considering user trust. Our findings suggest LLM assistants can be improved by: 1) training or prompting LLMs to estimate user trust based on conversational history, 2) training LLMs to adaptively decide when and how to insert trust-adaptive interventions in their generations to support reliable user decision-making.
Further, while free-text LLM explanations have been widely studied (Rajani et al., 2019; Wiegreffe et al., 2022), it is unclear whether they are actually useful to users (Joshi et al., 2023; Si et al., 2024; Hashemi Chaleshtori et al., 2024). Our work adds to this discourse by finding that LLM-generated explanations are useful for mitigating inappropriate reliance when user trust is low or high.
Applying trust-adaptive interventions.
Is all inappropriate reliance equally bad?
Should AI assistants be selectively transparent?
The research community has historically regarded transparency about AI behavior and reasoning through model explanations as essential for responsible and accountable deployment of AI systems (Schmidt et al., 2020; Liao and Vaughan, 2023; Felzmann et al., 2020). However, many previous works have found that explanations can often be at odds with optimal decision-making performance by causing users to over-rely on AI assistance (Zhang et al., 2020; Bansal et al., 2021). There is therefore a need to design explainable AI systems that balance transparency with appropriate reliance (de Souza Filho et al., 2024). However, many previous works have found that explanations can often be at odds with optimal decision-making performance by causing users to over-rely on AI assistance (Zhang et al., 2020; Bansal et al., 2021). There is therefore a need to design explainable AI systems that balance transparency with the need to encourage appropriate reliance (de Souza Filho et al., 2024). To this end, the community has investigated the utility of selectively transparent systems (Muralidhar, 2024; Muralidhar et al., 2025; François et al., 2022), where explanation details may be disclosed progressively (Muralidhar, 2024) or when demanded by users (Muralidhar et al., 2025). Our work extends this line of research by conditioning when and how to be transparent about AI reasoning on users’ trust levels. We find that using trust as a cue to modulate AI transparency improves users’ decision-making and reduces inappropriate reliance. However, selective transparency may not be appropriate for all contexts; we urge AI system deployers to tailor the degree of AI transparency based on the specific deployment context and the transparency-performance trade-off for their application.
Limitations.
The majority of our user studies were performed using a simulated AI assistant, rather than a real system such as a Large Language Model. Similar to Buçinca et al. (2021); Dhuliawala et al. (2023), we use a simulated AI for the controllability of AI accuracy and confidence calibration. However, a simulated AI may be unrealistic compared to real AI systems that people use. For instance, since the answer correctness is decided independently for each problem, our AI assistant may provide contradicting predictions for near-identical problems.
Furthermore, the behavior of Prolific participants interacting with an AI in a user study may differ from users of a live system in a real-world scenario (McGrath, 1995), particularly due to misalignment of motivations (Deci et al., 1999). We test for the effect of two confounding variables: trust reporting and user experience with AI (Appendix D), finding that they do not affect user reliance behavior, but there may be other confounding variables we have not yet considered.
Finally, our findings would be more reliable if we could obtain data from more participants, more tasks and more interventions, made difficult due to financial restrictions and the dearth of professional doctors on the Prolific platform.
Ethical Considerations.
13. Conclusion and Future Directions
Our findings present an initial exploration into adapting AI behaviors based on user trust levels. We adopted a simple thresholding criterion for deciding when to intervene, but more sophisticated criteria that also account for user and AI confidence may have potential. Further, rather than looking at whether user trust is too high or low, we can consider whether the trust is calibrated with the assistant’s trustworthiness. For example, high user trust may not be as undesirable when the AI is significantly more accurate than the user on the task. We hope our findings inspire the community to more closely consider the effect of user trust in user-AI interactions, and to invest more resources into the study of modeling user trust.
Acknowledgements.
We thank Rishi Vanukuru, Gregory Yauney, Sayan Ghosh and Jesse Zhang for their invaluable discussions and brainstorming, as well as Malihe Alikhani, Heather Culbertson, Suvodip Dey, Vardhan Dongre, Mert İnan, Sunnie Kim, Chaitanya Malaviya, Ana Marasović, Weiyan Shi, Anthony Sicilia, Swabha Swayamdipta, Dilek Hakkani-Tür and Gökhan Tür for their valuable feedback. We also thank Dhuliawala et al. (2023) for open-sourcing their user study interface, and in particular Vilém Zouhar for his assistance with setting up the user study and data logging. We are also very grateful to Aisha Baptiste for swiftly adding money to our Prolific account whenever required. Tejas Srinivasan was supported by the Amazon ML Fellowship from the USC-Amazon Center for Secure and Trusted Machine Learning. This work was supported in part by a grant from the DARPA Friction and Accountability in Conversational Transactions (FACT) AI Exploration (Award HR00112490376) and in part by a grant from USC-Capital One Center for Responsible AI and Decision Making in Finance (CREDIF) to the first author. The opinions and findings expressed in this work do not necessarily reflect those of USC, Amazon, Capital One, or DARPA.
Appendix A Task Details
ARC Task.
This task consists of answering science questions. Questions are sourced from the ARC dataset (Clark et al., 2018), which consists of more than 7000 grade-school science multiple-choice questions written for examinations. The authors manually reviewed questions from this dataset and selected questions that were challenging (i.e. the correct answer is not immediately obvious, and at least one option was not obviously incorrect) but still understandable (did not contain any scientific jargon that laypeople may not be familiar with). All questions in the original dataset had four options, but the author only selected the correct answer and the most plausible incorrect option for the decision-making problem. The final filtered set consisted of 39 questions. For each of the 10 problem sequences that users solve, we sample 30 of the 39 questions without repetition.
Diagnosis Task.
This task consists of diagnosing patients based on patient symptoms. Patient symptoms are sourced from the DDXPlus dataset (Fansi Tchango et al., 2022), which contains 1.3 million synthetic patients with a differential diagnosis. The symptoms are presented as either binary, categorical or continuous variables, but each symptom has a corresponding patient intake question and answer in English, which we translate into a descriptive third-person statement using GPT-4o.. For example, the intake question “Do you feel pain somewhere?” and patient response “Knee (R)” is translated into “The patient feels pain in their right knee.”. We filter down to questions with only 10–15 intake responses so that users do not need to spend a long time understanding the problem. To convert the task into a multiple-choice problem, we select the top three negative conditions from the differential diagnosis as the incorrect options. Our final set of problems includes 55 cases corresponding to eleven different conditions, which we use to sample 10 sequences of 30 problems each.
Appendix B User Study Details
Figure 11. Macro-aggregation results for mitigating under- and over-reliance. $n$ represents the number of users who have at least 3 interactions meeting the required criteria.
Macro-aggregation results for mitigating under- and over-reliance.
We present additional details about the user studies.
User Payment.
As mentioned earlier, users in the ARC task are paid a base payment of \$1.0, with an incentive of \$0.10 per correct answer. Users achieved 65–75% accuracy on the task, which translates to a bonus of $\approx\mathdollar 2.0$ per user, or \$3.0 total payment. The tasks took a median time of 15 minutes to complete, which translates to a pay rate of \$12 per hour.
User Demographics.
Instructions.
Users were informed that they were participating in a research study. Figure 10 contains screenshots of the task instructions and payment details presented to users. To maintain user attention, participants were instructed not to change their browser tab during the study, and a pop-up notification would send a warning whenever they did change their tab. To further ensure that users read the questions and choices closely before making their initial decision, the buttons for selecting the initial decision are only made active after 10 seconds.
Appendix C Macro-Aggregation Results
In Figure 11, we observe trends similar to those observed in Section 5.3. However, when looking at the low/high trust subsets, most conditions have fewer than 10 users for calculating Under-Reliance and Over-Reliance, and fewer than 5 users for Total Inappropriate Reliance.
Appendix D Testing for Confounding Variables